MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 19, 2010.
Published byModified over 5 years ago
Presentation on theme: "MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 19, 2010."— Presentation transcript:
MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 19, 2010
Commissioning Needs-Soler2 Commissioning Needs Decay Solenoid: —Summary of work on the Linde refrigerator last week —Linde sent their engineer, Matthias Hoberich, and was available to work on Linde fridge between Wed-Fri 14-16 April). —HPC received shaft seal Mon 12 April and sent engineer on Wed 14 April. —HPC tried to extract seal on Wed 14 April but it was lodged inside and would not move. —Returned on Thurs 15 April with an extraction tool. He dismantled the compressor and tried to extract the shaft seal but the tool was the wrong size. —Finally the HPC engineer decided to take the compressor back to HPC, to have it removed on the bench. —They will return tomorrow Tuesday 20 April with a new compressor since it is suspected that the fact that the seal was stuck on compressor shaft might have been cause of the problem. —It made no sense for Matthias to remain at RAL, so he returned by train to Frankfurt last Friday, where he had another job. —It is possible that Matthias could return to RAL this Friday, but travel disruptions due to the volcanic ash problem might make this unlikely. —At the moment it is uncertain whether this will be finished this week. —Even when finished, it will take ~10 days to cool down magnet.
April 19, 2010Commissioning Needs-Soler3 Commissioning Needs Target — ISIS target working well Detectors —Luminosity Monitor o Dedicated machine physics run last Friday 16 April to determine dead time —Tracker o Alan Bross at RAL last week (and this week!) to fix problem with leaks. o Sumitomo engineer replaced heads and now ready to cool down o Will debug DAQ before cooling down —CKOV,TOF0, TOF1, TOF2, KL: all these detectors are powered up and ready o TOF0,1,2: need calibration o Need beam configuration from Marco to blow up beam to populate all TOF area —Beam Line Monitors: o These were running on Friday for the machine development run but could not read them out (need DAQ expert). Beam Magnets —All beam magnets (except DS) worked well last Friday
April 19, 2010Commissioning Needs-Soler4 Preparation for run Machine development period: 12-19 April —We ran one shift (18:00-1:30) on Friday 16 April for Luminosity Monitor rate vs gate width studies (change LM gate from 10-40 ns and beamloss 0.5-2.0V) o Unfortunately, beam was unstable and could not finish programme of work o Need to reschedule this run for users run, probably on 20 April. Programme ISIS run: 20 April – 28 May —While decay solenoid not operational (20 April – first or second week May?) o Reschedule luminosity monitor gate studies for Tue 20 April 18:00-24:00 o Perform new TOF0,1,2 calibration with blown up defocused beam to try and cover all counters – would need beam parameters from Marco o Request from TOF experts (Maurizio, Ludovico) to have shorter runs since DAQ becomes unsynchronised and most data at the end of runs is unusable – proposal to have runs lasting only ~10-15 mins o Proton absorber tests – move different thickness of absorber to compare particle rate vs absorber thickness —While decay solenoid operational (First or second week May-26 May): o Perform target scan: delay target vs depth (ie acceleration) to try and flatten spill and reduce pileup and beam overlaps. o Work our way through Matrix of emittance (4 , 6 , 10 ) vs momentum (160, 200, 240 MeV/c) with all detectors on. o Statistics for each point? Maybe ~10 5 muons
April 19, 2010Commissioning Needs-Soler5 Preparation for run Proposal for shifts 20 April-28May: —Inefficient to set up runs (~ 2 hours preparation) if work being carried out in hall —Can only work in evening (from 18:00 onwards) —Proposal to have a few evening shifts but to have bulk of running over weekends between 18:00 Friday to 7:00 Monday —We would need to have at least two shifts each day with two people on shift (for safety). —I propose to finish luminosity monitor gate running this Tuesday (or Wednesday if programme not complete or beam not good) from 18:00 onwards —Then carry out remaining programme of work in weekend periods. —Have more shorter runs to avoid DAQ problems